OpenStudy (anonymous):

sin((1/2)arccos(3/5))

zepdrix (zepdrix):

Hey there! Welcome to OpenStudy.\[\Large\rm \sin\left[\frac{1}{2}\arccos\frac{3}{5}\right]\]Hmm ok let's see how we should approach this....

zepdrix (zepdrix):